OpenStudy (anonymous):

MEDAL AND FAN! Forty is 25% of what number?

OpenStudy (asnaseer):

Here again you can make use of the fact that percent means "out of 100". So, 25% means \(\frac{25}{100}\) So now you need to work out some number such that 25% of it will equal 40. If we call this unknown number \(x\), then this translates to:\[40=\frac{25}{100}\times x\]

OpenStudy (asnaseer):

You now need to use this equation to work out what \(x\) is.

OpenStudy (anonymous):

Sorry! Thought it was a simple question so I just gave answer, and I'm new :P

OpenStudy (anonymous):

i got help already

OpenStudy (anonymous):

160

OpenStudy (asnaseer):

@SuperNinjaPower - np - hope you understood the method as well :)